I would like you to help me on solving this problem.
The perimeter of a rectangle is 32cm. Find it's dimension if the length is 4cm longer than it's width.

Answer by ankor@dixie-net.com(22740)

The perimeter of a rectangle is 32cm.
2L + 2W = 32
Simplify, divide thru by 2, results:
L + W = 16
:
the length is 4cm longer than it's width.
L = (W+4)
:
Find it's dimension
L + W = 16
replace L with (W+4)
(W+4) + W = 16
2W = 16-4
2W = 12
W = 12/2
W = 6 cm is the width
:
You can find the Length, check it in the perimeter equation